For multimode step-index fiber with core index of 1.68 and radius of 35 um. The indexdifference is 3.75%. For a wavelength of 1150 nm, the optical power propagating incladding is what percentage of total power?(a) 2.1%(b) 2.5%(c) 1.8%(d) 1.4%​

1
Expert's answer
2021-03-08T07:27:55-0500

Given,

Core index =1.68

Radius "=35\mu m"

Index difference "=3.75\%"

wavelength "(\lambda) =1150nm = 1150\times 10^{-9}m"

"0.0375=1.68 \times I_1"


"I_1=\frac{0.0375}{1.68}= 0.02232"


Index "=I_o-I_1"

"=1.68-0.02232"

"=1.65"

"P=\frac{1}{2} I_o(\pi W^2)"


"P=\frac{1}{2} \times 1.65 \times 3.14 \times 0.1150"

"=0.0297"

"P=2.5\%"